OpenStudy (anonymous):

using the shell method to find volume if you are rotating about the axis x=-5, is the radius x+5 or x-5?

OpenStudy (jebonna):

x saiken200

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

the radius is only x if it's rotating around the y axis itself. it's not. it's rotating around the axis x=-5. so it's -5 away from the y axis.

OpenStudy (anonymous):

for the love of god will someone please help me!!!!!

OpenStudy (anonymous):

i have tried this damn problem 6 different ways and cannot get the right damn answer!!!!!!!Q~!Q!@w'x

OpenStudy (jebonna):

its x + 5

OpenStudy (anonymous):

ok, answer's still wrong.... The region bounded by y=x8 and y=sin(πx/2) is rotated about the line x=−5.

OpenStudy (anonymous):

\[2\pi \int\limits_{0}^{1}(x+5)(x^8-\sin(\pi x/2)\]

OpenStudy (anonymous):

i integrate that, plug in 1-0 and get it wrong....this assignment is due in a few hours! i NEED HELP!!!

OpenStudy (jebonna):

i am an illegal immegrant.. my nicknake is 2 bags.. i found a little bit of sheeeeeet on my ppe.. i flicked it off.. like i didnt know it was there.. im sorry...

OpenStudy (anonymous):

how is that supposed to be an answer? i seriously need help.

OpenStudy (amistre64):

r = k - f(x), if you are trying to assess a radius otherwise you can just shift the setup to the y axis .... by f(x-c)

OpenStudy (phi):

@saiken2009 you set up the problem as *** \( 2\pi \int\limits_{0}^{1}(x+5)(x^8-\sin(\pi x/2)\) ***** but it is better to write it as \[ 2\pi \int\limits_{0}^{1}(x+5)(\sin(\pi x/2)- x^8) \] because the sin is the "top" curve

OpenStudy (phi):

if you multiply it out you get \[ 2\pi \int\limits_{0}^{1}(x+5)(\sin(\pi x/2)- x^8) \ dx \\ 2\pi \int\limits_{0}^{1} x \sin\left(\frac{\pi x}{2}\right)+ 5 \sin\left(\frac{\pi x}{2}\right)-x^9 -5x^8 \ dx \]

OpenStudy (phi):

the first integral can be done using "integration by parts" the 2nd integral is straight forward: \[ 2\pi \int\limits 5 \sin\left(\frac{\pi x}{2}\right)= 2\pi \cdot 5 \cdot \frac{2}{\pi} \cdot -1 \cdot \cos\left(\frac{\pi x}{2}\right) = -20 \cos\left(\frac{\pi x}{2}\right) \] as are the last 2: \[ 2\pi \int -x^9= -\frac{2 \pi}{10} x^{10}\\ 2\pi \int -5x^8= -\frac{10 \pi}{9} x^{9} \]